The equation of the tangent plane to h(x,y)=yln(x) at the point (1,5) is
Select one:
a. z=1/5x+1/5y

b. z=x+y


c. z=1/4x+1/4y


d. z=4(x−1)


e. z=5(x+1)


f. z=5(x−1)
1
Expert's answer
2020-05-27T16:29:58-0400

"Since \ h(x,y)=yln(x), we \ get \\
h_x(x,y)=\frac{y}{x} \ and \\
h_y(x,y)=\ln(x).\\
at (1,5) , we \ obtain\\

h_x(1,5)= 5\ and \\
h_y(1,5)=\ln(1)=0.\\
As \ at (1,5)
 \text{the equation of }\\ \text{the tangent plane to }h(x,y) is \\
z=h_x (1,5)(x-1)+h_y(1,5)(y-5)\\
So, we \ get \\
z=5(x-1).\\
\text{This implies that the correct answer is }\\
f."